# Multi-Level Damage-Aware Graph Learning for Resilient UAV Swarm Networks

### v2

@@ VI Conclusion
In this paper, we study the CNS issue of the USNET under massive damage scenarios. To cope with the challenges of utilizing graph-learning approaches to generate optimal recovery trajectories, we propose a novel ML-DAGL algorithm. To adapt to diverse damage cases with varying scales, we introduce the MBDA module to form a sequence of mDAGs with different ranges of receptive fields. The mDAG links only remaining and damaged nodes to ensure a more even degree distribution for mitigating over-aggregation and utilizes multi-hop dilation to establish more links for sparse topology enhancement. We then develop the DGCN to generate the optimal recovery trajectories based on the mDAG sequence, which can ensure convergence under massive damage cases through B-GCO and residual connection. Statistical results demonstrated that the proposed algorithm can guarantee the connectivity restoration under varying swarm and damage scales, along with significant improvements in both recovery time and topology degree distribution.
@@ Appendix A Proof of Proposition 1
[…]
@@
$=\mathop{\rm max}\limits_{1\leq i\leq N}\{|1\!-\!\epsilon d_{i}^{k}|\!+\!\mathop{\sum}\limits_{j=1}^{N}\epsilon a_{ij}^{k}\}\|\bm{X}_{gco}^{a}\!-\!\bm{X}_{gco}^{b}\|_{\infty}$
@@
[…]
@@ Appendix B Proof of Proposition 2
The DGCN consists of an input normalization layer, $Q$ B-GCO layers, $Q-1$ nonlinear activation layers, and an output upscale layer. Compared to other operations, the graph convolution within the B-GCO layers takes the most computational complexity. For each B-GCO layer, the GCO applies matrix multiplications for two times in (24). The computational complexity of a single B-GCO layer is the sum of the complexities of matrix multiplications.

### v3

@@ V-E3 Study of DGCN
The proposed DGCN includes two key components designed to solve CNS problems: the B-GCO and a residual connection mechanism. The average recovery time from ablation experiments on DGCN is shown in Fig.9, where the B-GCO is replaced by the original GCO from [25], and the residual connection is optionally removed. The results show that omitting the residual connection leads to increased recovery time, indicating reduced convolutional efficiency on sparse topologies. This confirms the effectiveness of the residual connection in handling large-scale damage. In parallel, B-GCO outperforms the original GCO by an average margin of 22.4%. This improvement stems from their different normalization strategies: the original GCO applies degree normalization at the node level, while B-GCO performs at the graph level to preserve global degree-weight information. Consequently, DGCN can more effectively leverage the original network structure to address CNS challenges.
V-F Scalability analysis
TABLE I: Statistical results of resilient algorithms under half-damaged scenarios with different swarm scales.
  Algorithm
$N=20$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
$N=50$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
$N=100$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
center-fly $\textbf{1.00}\,/\,1.95\,/\,1.35\,/\,4.52\,/\,9$ $\textbf{1.00}\,/\,8.81\,/\,6.13\,/\,10.25\,/\,24$ $\textbf{1.00}\,/\,11.02\,/\,7.53\,/\,11.73\,/\,49$
[…]
SIDR $0.56\,/\,8.07\,/\,7.11\,/\,\textbf{3.56}\,/\,\textbf{8}$ $0.70\,/\,9.90\,/\,10.11\,/\,\textbf{3.82}\,/\,\textbf{11}$ $0.52\,/\,25.93\,/\,12.02\,/\,\textbf{3.97}\,/\,\textbf{11}$
GCN $\textbf{1.00}\,/\,5.09\,/\,2.81\,/\,5.00\,/\,9$ 0.98/ 12.22/ 5.00/ 7.38/ 200.98\,/\,12.22\,/\,5.00\,/\,7.38\,/\,20 0.90/ 20.52/ 9.91/ 10.73/ 340.90\,/\,20.52\,/\,9.91\,/\,10.73\,/\,34
GAT $\textbf{1.00}\,/\,5.02\,/\,3.63\,/\,4.88\,/\,9$ $\textbf{1.00}\,/\,10.28\,/\,4.61\,/\,7.74\,/\,20$ $\textbf{1.00}\,/\,14.83\,/\,5.70\,/\,9.42\,/\,26$
CR-MGC $\textbf{1.00}\,/\,2.80\,/\,2.59\,/\,4.64\,/\,9$ 0.98/ 9.00/ 5.53/ 8.44/ 230.98\,/\,9.00\,/\,5.53\,/\,8.44\,/\,23 $\textbf{1.00}\,/\,11.14\,/\,5.89\,/\,9.00\,/\,30$
DEMD $\textbf{1.00}\,/\,1.92\,/\,1.39\,/\,4.49\,/\,9$ $\textbf{1.00}\,/\,7.78\,/\,5.00\,/\,8.74\,/\,23$ $\textbf{1.00}\,/\,10.54\,/\,4.91\,/\,8.22\,/\,35$
ML-DAGL $\textbf{1.00}\,/\,\textbf{1.74}\,/\,\textbf{1.19}\,/\,3.97\,/\,9$ $\textbf{1.00}\,/\,\textbf{3.34}\,/\,\textbf{2.38}\,/\,4.24\,/\,12$ $\textbf{1.00}\,/\,\textbf{3.72}\,/\,\textbf{1.71}\,/\,4.53\,/\,\textbf{11}$
  Algorithm
$N=200$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
$N=500$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
$N=1000$
$R_{c}\,/\,E(T_{rc})\,/\,\sigma(T_{rc})\,/\,E(d)\,/\,max(d)$
center-fly 0.98/ 22.90/ 11.17/ 26.54/ 980.98\,/\,22.90\,/\,11.17\,/\,26.54\,/\,98 0.88/ 46.70/ 18.81/ 71.04/ 2380.88\,/\,46.70\,/\,18.81\,/\,71.04\,/\,238 0.50/ 90.12/ 26.16/ 260.78/ 4630.50\,/\,90.12\,/\,26.16\,/\,260.78\,/\,463
HERO $^{\dagger}\,/\,-\,/\,-\,/\,-\,/\,- 0////0^{\dagger}\,$/\,-\,/\,-\,/\,-\,/\,- $0^{\dagger}\,/\,-\,/\,-\,/\,-\,/\,$-
SIDR $0.36\,/\,34.81\,/\,20.43\,/\,\textbf{4.53}\,/\,16$ $0.32\,/\,69.94\,/\,14.81\,/\,\textbf{4.19}\,/\,\textbf{13}$ $0.22\,/\,90.64\,/\,41.40\,/\,\textbf{4.39}\,/\,\textbf{15}$
GCN 0.86/ 30.12/ 12.72/ 12.14/ 630.86\,/\,30.12\,/\,12.72\,/\,12.14\,/\,63 0.86/ 58.52/ 14.52/ 56.72/ 2140.86\,/\,58.52\,/\,14.52\,/\,56.72\,/\,214 $0.08^{\dagger}\,/\,-\,/\,-\,/\,-\,/\,$-
GAT $\textbf{1.00}\,/\,24.48\,/\,9.64\,/\,13.32\,/\,65$ 0.94/ 56.90/ 12.98/ 21.81/ 960.94\,/\,56.90\,/\,12.98\,/\,21.81\,/\,96 0.80/ 95.83/ 13.50/ 38.12/ 2190.80\,/\,95.83\,/\,13.50\,/\,38.12\,/\,219
CR-MGC $\textbf{1.00}\,/\,21.28\,/\,8.68\,/\,17.66\,/\,84$ 0.76/ 63.84/ 14.34/ 29.39/ 1080.76\,/\,63.84\,/\,14.34\,/\,29.39\,/\,108 0.34/ 104.13/ 15.62/ 33.38/ 1880.34\,/\,104.13\,/\,15.62\,/\,33.38\,/\,188
DEMD $\textbf{1.00}\,/\,21.40\,/\,7.32\,/\,13.37\,/\,53$ $\textbf{1.00}\,/\,45.40\,/\,8.72\,/\,18.14\,/\,67$ 0.68/ 86.89/ 22.79/ 25.77/ 1400.68\,/\,86.89\,/\,22.79\,/\,25.77\,/\,140
ML-DAGL $\textbf{1.00}\,/\,\textbf{5.24}\,/\,\textbf{2.10}\,/\,5.24\,/\,\textbf{15}$ $\textbf{1.00}\,/\,\textbf{18.10}\,/\,\textbf{7.69}\,/\,7.46\,/\,30$ $\textbf{1.00}\,/\,\textbf{32.58}\,/\,\textbf{6.18}\,/\,8.53\,/\,39$
To evaluate the scalability of the proposed algorithm for varying swarm scales, we set the number of UAVs as $N=50,100,200,500,1000$. The damage ratio is set to $p=0.5$, and each simulation of different $N$ is repeated 50 times randomly for the statistical results. Simulation results for resilience algorithms are summarized in TABLE I, including the convergent ratio $R_{c}$, mean recovery time $E(T_{rc})$, variance of recovery time $\sigma(T_{rc})$, average node degree $E(d)$, and maximum node degree $\max(d)$.
The results show that ML-DAGL consistently maintains $R_{c}=1$ with full connectivity restoration across different swarm sizes, demonstrating strong scalability. In contrast, other algorithms exhibit declining $R_{c}$ as the swarm size increases, especially with poor performance when $N$ reaches 1000. Meanwhile, regarding recovery efficiency, ML-DAGL achieves state-of-the-art results in both the mean and variance of recovery time, enabling more efficient and stable recovery solutions. For metrics on degree distribution, the ML-DAGL yields the second lowest values for both $E(d)$ and $\max(d)$, only inferior to the distributed algorithm SIDR. In summary, the proposed algorithm has excellent scalability.
@@ VI Conclusion
In this paper, we studied the CNS issue of the USNET under massive damage scenarios. To cope with the challenges of utilizing graph-learning approaches to generate optimal recovery trajectories, we proposed a novel ML-DAGL algorithm. To adapt to diverse damage cases with varying scales, we introduced the MBDA module to form a sequence of mDAGs with different ranges of receptive fields. Each mDAG maintains the edges between remaining and destroyed nodes within the k-hop range, establishing dilated topologies with damage attention. We then developed the DGCN to generate the optimal recovery trajectories based on the mDAG sequence, utilizing the bipartite GCO and residual connection to enhance the feature mining and aggregation. Additionally, we theoretically proved the convergence of the proposed algorithm and analyzed the upper bound of recovery time and computational complexity. The proposed algorithm is feasible for practical deployment with limited memory requirement. Simulations demonstrated that ML-DAGL algorithm have excellent scalability, along with significant improvements in convergence, recovery time, and topology distribution.
@@ Appendix A Proof of Proposition 1
Since the infinity norm of matrix $\|\cdot\|\infty$ has the sub-multiplicative property, we can get
@@
$=\mathop{\rm max}\limits_{1\leq i\leq N}\{|1\!-\!\epsilon d_{i}^{k}|\!+\!\mathop{\sum}\limits_{j=1}^{N}\epsilon a_{ij}^{k}\}\|\bm{X}_{gco}^{a}\!-\!\bm{X}_{gco}^{b}\|_{\infty}$ .
@@
Given the proof of the sub-multiplicative property of $\|\cdot\|\infty$ as
@@ Appendix B Proof of Proposition 2
The DGCN consists of an input normalization layer, $Q$ B-GCO layers, Q1Q-1 nonlinear activation layers, and an output upscale layer. Compared to other operations, the graph convolution within the B-GCO layers takes the most computational complexity. For each B-GCO layer, the GCO applies matrix multiplications for two times in (25). The computational complexity of a single B-GCO layer is the sum of the complexities of matrix multiplications.